# MultiSensor-Home: A Wide-area Multi-modal Multi-view Dataset for Action Recognition and Transformer-based Sensor Fusion

### v1

@@ V-B Quantitative Results
Tables IV(a) and IV(b) present the results of MultiTSF compared with other methods on the MultiSensor-Home and MM-Office datasets, respectively.
On the MultiSensor-Home dataset (Table IV(a)), MultiTSF achieves the best performance across all metrics and settings. Specifically, MultiTSF significantly outperforms MultiTrans and MultiASL in the multi-modal setting, achieving 64.48% of mAPC and 87.91% of mAPS in the sequence-level setting (i.e., where only video sequence-level labels are available) and 76.12% of mAPC and 91.45% of mAPS in the frame-level setting (i.e., where frame-level labels are available). In the uni-modal setting, MultiTSF also surpasses TimeSformer, ViViT, and X-CLIP, demonstrating its robustness even when audio features are excluded.
On the MM-Office dataset (Table IV(b)), MultiTSF maintains its superior performance in the sequence-level setting. In the uni-modal visual configuration, MultiTSF achieves the highest mAPC of 81.71% and mAPS of 91.23%, outperforming all baselines, including MultiASL and MultiTrans. In the multi-modal setting, MultiTSF achieves competitive results with 93.03% of mAPS, surpassing MultiASL and demonstrating its ability to effectively integrate audio and visual modalities.
@@ V-C Qualitative Results
Figure 5 illustrates the attention heatmaps from the Shared Visual Encoder on the MultiSensor-Home dataset, demonstrating its ability to identify action-relevant regions across both spatial and temporal dimensions. Figure 5(a) shows multi-view inputs alongside the attention heatmaps, revealing that the model effectively focuses on key objects such as lamps, curtains, desks, and human presence, which are critical for action recognition. Figure 5(b) visualizes the temporal attention maps, demonstrating how the model tracks significant changes in regions of interest over time, such as human movement and interaction with objects.
@@
Additionally, Figure 6 visualizes the attention scores from the Transformer-based Sensor Fusion across multiple views on the MultiSensor-Home dataset, highlighting the ability of the model to dynamically assign importance to different views based on their relevance to specific actions, such as “Turn On Lamp” and “Sit Down”. The attention scores demonstrate the model’s focus on the most informative perspectives (i.e., View 4 and View 5), further validating the effectiveness of the proposed fusion mechanism.
@@ V-D Ablation Studies
To evaluate the effectiveness of individual components in MultiTSF, we conducted a series of ablation studies on the MultiSensor-Home dataset.
@@ V-D1 Effectiveness of Loss Components
Table IV shows the contributions of each loss component. The results demonstrate that the full model (row 1) achieves the highest performance in both the uni-modal and multi-modal settings. Notably, excluding $\mathcal{L_{\text{F}}}$ (row 2) or $\mathcal{L_{\text{S}}}$ (row 3) leads to noticeable performance degradation, emphasizing the importance of frame-level and sequence-level supervision. The removal of $\mathcal{L_{\text{H}}}$ (rows 4, 5, 6) results in a significant performance drop, particularly in the multi-modal setting, underscoring its essential role in guiding the model to focus on human-centric regions for effective action recognition. Additionally, replacing $\mathcal{L_{\text{F}}}$ and $\mathcal{L_{\text{S}}}$ with Cross Entropy loss (row 7) further reduces accuracy, highlighting the superiority of Two-Way loss in handling class imbalance effectively.
V-D2 Device Fusion Strategies
We compare the proposed Transformer-based Sensor Fusion with other widely used fusion strategies, including max pooling, mean pooling, and concatenation. Table V shows that the Transformer-based approach achieves the best results across both uni-modal and multi-modal settings. While mean pooling performs competitively in certain metrics, its inability to adaptively model inter-device relationships limits its overall effectiveness. Similarly, concatenation captures features from multiple devices but lacks the capability to prioritize relevant views dynamically. In contrast, the proposed Transformer-based fusion excels by effectively modeling inter-device dependencies and assigning importance to the most relevant views, resulting in superior performance across all metrics.
@@ V-D2 Device Fusion Strategies
No. $\mathcal{L_{\text{H}}}$ $\mathcal{L_{\text{S}}}$ $\mathcal{L_{\text{F}}}$ Uni-modal (Visual) Multi-modal
@@
Change $\mathcal{L_{\text{V}}}$ and $\mathcal{L_{\text{F}}}$ to Cross Entropy loss
@@
TABLE V: Effectiveness of device fusion strategies. The best and second-best results are highlighted in bold and underlined text, respectively.
Device Fusion Strategies Uni-modal (Visual) Multi-modal
@@ VI Conclusion
In this study, we introduced the Multi-modal Multi-view Transformer-based Sensor Fusion (MultiTSF) method and the MultiSensor-Home dataset to address challenges in multi-modal multi-view action recognition. The MultiSensor-Home dataset provides a robust benchmark with untrimmed, multi-view sensors and detailed frame-level annotations. The proposed MultiTSF method demonstrated state-of-the-art performance on the proposed MultiSensor-Home dataset and MM-Office dataset [35] by effectively fusing audio and visual modalities through a Transformer-based mechanism and leveraging human detection for enhanced spatial learning. The results highlight significant improvements in mAPC (macro-averaged metric) and mAPS (micro-averaged metric), establishing MultiTSF as a competitive solution for real-world applications.
@@ Acknowledgment
This work was partly supported by JSPS KAKENHI JP21H03519 and JP24H00733.

### v3

@@ V-B Quantitative Results
Tables IV(a) and IV(b) present the results of the proposed MultiTSF compared with other methods on the MultiSensor-Home and MM-Office [35] datasets, respectively.
On the MultiSensor-Home dataset (Table IV(a)), MultiTSF achieved the best performance across all metrics and settings. Specifically, MultiTSF significantly outperformed MultiTrans [35] and MultiASL [18] in the multi-modal setting, achieving 64.48% of mAPC and 87.91% of mAPS in the sequence-level setting (i.e., where only video sequence-level labels were available) and 76.12% of mAPC and 91.45% of mAPS in the frame-level setting (i.e., where frame-level labels were available). In the uni-modal setting, MultiTSF also surpassed TimeSformer [3], ViViT [1], and X-CLIP [16], demonstrating its robustness even when audio features were excluded.
On the MM-Office dataset (Table IV(b)), MultiTSF maintained its superior performance in the sequence-level setting. In the uni-modal visual configuration, MultiTSF achieved the highest mAPC of 81.71% and mAPS of 91.23%, outperforming all baselines, including MultiASL and MultiTrans. In the multi-modal setting, MultiTSF achieved competitive results with 93.03% of mAPS, surpassing MultiASL and demonstrating its ability to effectively integrate audio and visual modalities.
@@ V-C Qualitative Results
Figure 5 illustrates the attention heatmaps from the Shared Visual Encoder on the MultiSensor-Home dataset, demonstrating its ability to identify action-relevant regions across both spatial and temporal dimensions. Figure 5(a) shows multi-view inputs alongside the attention heatmaps, revealing that the model effectively focused on key objects such as lamps, curtains, desks, and human presence, which are critical for action recognition. Figure 5(b) visualizes the temporal attention maps, demonstrating how the model tracked significant changes in regions-of-interest over time, such as human movement and interaction with objects.
@@
Additionally, Figure 6 visualizes the attention scores from the Transformer-based Sensor Fusion across multiple views on the MultiSensor-Home dataset, highlighting the ability of the model to dynamically assign importance to different views based on their relevance to specific actions, such as “TurnOnLamp” and “SitDown”. The attention scores demonstrate the model’s focus on the most informative perspectives (i.e., Views 4 and 5), further validating the effectiveness of the proposed fusion mechanism.
@@ V-D Ablation Studies
To evaluate the effectiveness of individual components in the proposed MultiTSF, we conduct a series of ablation studies on the MultiSensor-Home dataset.
@@ V-D1 Effectiveness of Loss Components
Table IV shows the contributions of each loss component. The results demonstrate that the full model (row 1) achieved the highest performance in both the uni-modal and multi-modal settings. Notably, excluding $\mathcal{L_{\text{F}}}$ (row 2) or $\mathcal{L_{\text{S}}}$ (row 3) lead to noticeable performance degradation, emphasizing the importance of frame-level and sequence-level supervision. The removal of $\mathcal{L_{\text{H}}}$ (rows 4, 5, and 6) resulted in a significant performance drop, particularly in the multi-modal setting, underscoring its essential role in guiding the model to focus on human-centric regions for effective action recognition. Additionally, replacing $\mathcal{L_{\text{F}}}$ and $\mathcal{L_{\text{S}}}$ with Cross Entropy loss (row 7) further reduced accuracy, highlighting the superiority of Two-Way loss in handling class imbalance effectively.
V-D2 Sensor Fusion Strategies
We compare the proposed Transformer-based Sensor Fusion with other widely used fusion strategies, including max pooling, mean pooling, and concatenation. Table V shows that the Transformer-based approach achieved the best results across both uni-modal and multi-modal settings. While mean pooling performed competitively in certain metrics, its inability to adaptively model inter-device relationships limited its overall effectiveness. Similarly, concatenation captured features from multiple devices but lacked the capability to prioritize relevant views dynamically. In contrast, the proposed Transformer-based fusion report better accuracy by effectively modeling inter-device dependencies and assigning importance to the most relevant views, resulting in superior performance across all metrics.
@@ V-D2 Sensor Fusion Strategies
Row $\mathcal{L_{\text{H}}}$ $\mathcal{L_{\text{S}}}$ $\mathcal{L_{\text{F}}}$ Uni-modal (Visual) Multi-modal
@@
Change $\mathcal{L_{\text{S}}}$ and $\mathcal{L_{\text{F}}}$ to Cross Entropy loss
@@
TABLE V: Effectiveness of sensor fusion strategies. The best and second-best results are highlighted in bold and underlined text, respectively.
Sensor Fusion Strategies Uni-modal (Visual) Multi-modal
@@ VI Conclusion
In this study, we proposed the MultiSensor-Home dataset and the Multi-modal Multi-view Transformer-based Sensor Fusion (MultiTSF) method to address challenges in multi-modal multi-view action recognition. The MultiSensor-Home dataset provides a robust benchmark with untrimmed, multi-view sensors and detailed frame-level annotations. The proposed MultiTSF method demonstrated its superiority over state-of-the-art methods on the proposed MultiSensor-Home and MM-Office [35] datasets by effectively fusing audio and visual modalities through a Transformer-based mechanism and leveraging human detection for enhanced spatial learning. The results highlight significant improvements in mAPC (macro-averaged metric) and mAPS (micro-averaged metric), establishing MultiTSF as a competitive solution for real-world applications. In future work, we aim to optimize models for deployment and expand the MultiSensor-Home dataset to include more home environments.
@@ Acknowledgment
This work was partly supported by JSPS KAKENHI JP21H03519 and JP24H00733. The computation was carried out using the General Projects on supercomputer “Flow” at Information Technology Center, Nagoya University.
Ethical Impact Statement
In this paper, we introduced the MultiSensor-Home dataset and proposed the MultiTSF method for multi-modal multi-view action recognition. Below, we outline the ethical considerations and potential societal impacts of this study.
Data Collection and Privacy
The MultiSensor-Home dataset comprises video and audio recordings captured in an indoor environment using multiple synchronized sensors. The dataset was collected with explicit consent from participants, ensuring compliance with ethical guidelines on data privacy and protection. No personally identifiable information is included in the dataset, and all participants were informed about the purpose of data collection, usage, and potential applications. However, anonymizing the videos (e.g., through blurring or masking) is not feasible, as it would compromise the dataset’s utility and the validity of the research findings.
Bias and Fairness
The MultiSensor-Home dataset was designed to include diverse scenarios, such as variations in lighting conditions, clothing, and environmental settings, to ensure robustness across different contexts. However, potential biases may persist due to the data being collected from a single home environment. Future work should aim to expand the dataset to include a wider range of environmental conditions, improving fairness and generalizability.
Potential Misuse and Societal Impact
The proposed MultiTSF method aims to improve action recognition performance in various applications, including smart homes, assistive technologies, and human-computer interaction. However, like any computer vision technology, there is a risk of misuse, such as unauthorized surveillance or privacy-intrusive applications. To mitigate these risks, we encourage the responsible and ethical use of the proposed dataset and method in accordance with legal and ethical frameworks. The proposed dataset and method are intended solely for research purposes, and commercial usage should adhere to ethical AI principles.
Transparency and Reproducibility
To promote transparency and reproducibility, we provide detailed descriptions of the proposed MultiSensor-Home dataset and the MultiTSF method. The dataset and method implementation will be made publicly available under appropriate licensing to ensure responsible research practices and facilitate further advancements in the field.
Conclusion
In this paper, we contribute to the advancement of multi-modal multi-view action recognition while addressing ethical concerns related to privacy, bias, and potential misuse. We emphasize the importance of ethical AI practices and encourage future research to continue improving fairness, privacy safeguards, and responsible deployment of AI-driven action recognition technologies.